OpenStudy (anonymous):

solve -3v is less or equal to 12

OpenStudy (anonymous):

\[-3v \le 12\] simplify: divide by 3 \[-v \le 4\] swap sides since you want the positive of v \[-4 \le x\]

OpenStudy (anonymous):

\[v \ge -4\]

OpenStudy (anonymous):

Too used to write x as a variable :P